x^2-2/3x=0
I have to complete the square.

Answer by nerdybill(7384)

x^2-2/3x=0
take half of b, square it, and add to both sides (2/6)^2 = (1/3)^2 = 1/9:
x^2-2/3x+1/9 = 1/9
(x-1/3)^2 = 1/9
x-1/3 = sqrt(1/9)
x = sqrt(1/9) + 1/3
x = +-1/3 + 1/3
x = {0, 2/3}